Given f(x)=x^7 -x^5 -x^4 +2x +1 on interval [-1,1] show that there's at least one critical point on this interval.


1
Expert's answer
2022-02-09T17:36:13-0500
"f(x)=x^7 -x^5 -x^4 +2x +1"


The function "f(x)" is continuous and differentiable on "\R" as polynomial.

"f'(x)=7x^6-5x^4-4x^3+2""f'(0)=2>0"

"f'(0.8)=7(0.8)^6-5(0.8)^4-4(0.8)^3+2"

"=-0.260992<0"

Then by the Intermediate Value Theorem, there exists a number "c\in (0, 0.8)" such that "f'(c)=0."

Therefore there's at least one critical point on the interval"[-1,1]."